# Policy Contrastive Decoding for Robotic Foundation Models

### v4

@@ 2 Related Work
Contrastive Decoding. Recent studies in large vision-language models (VLMs) have highlighted the challenge of hallucination (Gunjal et al., 2024; Rawte et al., 2023; Liu et al., 2023b), where models generate inaccurate or misleading outputs not grounded in the input data. Contrastive Decoding (CD) (Favero et al., 2024; Chen et al., 2024; Leng et al., 2024) suppresses hallucinations by comparing output distributions from original and distorted input without altering the model’s weights. Instruction Contrastive Decoding (ICD) (Wang et al., 2024) is used to combat hallucinations in multimodal tasks by modulating the confidence in multimodal alignment of the model’s visual and textual inputs, enabling it to distinguish between hallucinated and relevant tokens. Visual Contrastive Decoding (VCD) (Park et al., 2025; Favero et al., 2024) aims to reduce object hallucinations by comparing outputs from original and distorted visual inputs. This approach provides a computationally efficient solution as it requires neither additional training nor external pre-trained models. Our proposed Policy Contrastive Decoding (PCD) approach in this work is mainly inspired by the success of VCD schemes. However, unlike those schemes that contrast output distributions from original and noise-distorted (Park et al., 2025) or image-removed (Favero et al., 2024) inputs to counteract LVLMs’ overreliance on language priors, our PCD contrasts action probability distributions derived from original and object-masked inputs, thereby redirecting the robot policy’s focus to object-relevant visual features and enhancing their generalization across different scenarios.
@@ 4 Experiments
Table 1: SIMPLER Performance. Task-specific objects in the initial observation are annotated by artificial Point and Box prompts or the automatic detection results of Grounding DINO (Liu et al., 2024b). The results are the success rate and the 95% confidence interval of 300 trials. As a plug-and-play approach, PCD consistently enhances the three policies by large margins over the 9 tasks.
@@ 5 Conclusion, Limitations and Future Work
Appendix A CAM Visualization of Failure Cases.
@@ C.2 Evaluation Tasks
Simulation
Close Drawer close top/middle/bottom drawer top/middle/bottom drawer 300
Move Near move X near Y
apple, orange, pepsi can, …
300
@@
Real-world
Pick Ball pick the ball ball 20
@@ Appendix E Computational Overhead
Appendix F LLMs Usage Statement

### v5

@@ 2 Related Work
Contrastive Decoding. Recent studies in large vision-language models (VLMs) have highlighted the challenge of hallucination (Gunjal et al., 2024; Rawte et al., 2023; Liu et al., 2023b; Zhang et al., 2025), where models generate inaccurate or misleading outputs not grounded in the input data. Contrastive Decoding (CD) (Favero et al., 2024; Chen et al., 2024; Leng et al., 2024) suppresses hallucinations by comparing output distributions from original and distorted input without altering the model’s weights. Instruction Contrastive Decoding (ICD) (Wang et al., 2024) is used to combat hallucinations in multimodal tasks by modulating the confidence in multimodal alignment of the model’s visual and textual inputs, enabling it to distinguish between hallucinated and relevant tokens. Visual Contrastive Decoding (VCD) (Park et al., 2025; Favero et al., 2024) aims to reduce object hallucinations by comparing outputs from original and distorted visual inputs. This approach provides a computationally efficient solution as it requires neither additional training nor external pre-trained models. Our proposed approach in this work is mainly inspired by the success of VCD schemes. However, unlike those schemes that contrast output distributions from original and noise-distorted (Park et al., 2025) or image-removed (Favero et al., 2024) inputs to counteract LVLMs’ overreliance on language priors, our PCD contrasts action probability distributions derived from original and object-masked inputs, thereby redirecting the robot policy’s focus to object-relevant visual features and enhancing their generalization across different scenarios.
@@ 4 Experiments
Table 1: SIMPLER Performance. Task-specific objects in the initial observation are annotated by artificial Point and Box prompts or the automatic detection results of GDINO (Liu et al., 2024b). The results are the success rate and the 95% confidence interval of 300 trials. As a plug-and-play approach, PCD consistently enhances the three policies by large margins over the 9 tasks.
@@ 5 Conclusion, Limitations and Future Work
Acknowledgement
This work is supported by grants from the National Natural Science Foundation of China (No.62506310, No.62425208, No.U22A2097, No.U23A20315, No.82441006), the China Postdoctoral Science Foundation (No.2025M781517), and the Natural Science Foundation of Sichuan Province (No.2026NSFSC1479).
Appendix A CAM Visualization of Failure Cases
@@ C.2 Evaluation Tasks
Simulation Close Drawer close top/middle/bottom drawer top/middle/bottom drawer 300
Move Near move X near Y apple, orange, pepsi can, … 300
@@
Real-world Pick Ball pick the ball ball 20
@@ Appendix E Computational Overhead
Appendix F PCD vs. Classifier-Free Guidance (CFG)
The proposed PCD algorithm shares similarity with the training-free approach classifier-free guidance (CFG) [Ho and Salimans, 2022], where the masked image serves as the unconditional inputs. However, the essential distinction lies in how and when the guidance is applied: CFG provides implicit guidance during the iterative decoding process. At each step of the diffusion, it steers the generation away from the unconditional prediction. This is an implicit probabilistic modeling process. In contrast, our PCD is an explicit post-hoc correction. The quantitative comparison shown in Table 6 shows that our PCD method significantly outperforms CFG across the nine SIMPLE tasks. However, the time cost of PCD is 1.29 times that of CFG. Therefore, harnessing the complementary strengths of PCD and CFG to further address spurious correlations remains a promising avenue for future research.
Table 6: Comparison of PCD and Classifier-Free Guidance (CFG) [Ho and Salimans, 2022].
Model $\pi_{0}$ +CFG +PCD
Average time cost for each inference step (seconds) 0.66 0.84 1.09
Average success rate over nine tasks (%) 63.9 62.7 68.1
Close Drawer 75.7 75.7 74.7
Move Near 67.3 62.7 66.7
Open Drawer 38.0 46.3 47.7
Pick Coke 84.0 81.3 84.3
Apple Drawer 17.0 13.0 26.0
Carrot Plate 58.0 59.0 67.7
Egg. Basket 86.0 87.3 83.7
Spoon Towel 80.7 78.3 86.3
Stack Cube 68.7 60.3 76.3
To further investigate CFG, we conduct a step-by-step analysis in Table 7, applying CFG after a certain step and running 100 trials for each setting. The results show a clear trend: the earlier the CFG intervention, the more severe the performance degradation. The primary reason appears to be that CFG’s influence on the noise space disrupts the semantic integrity of the model’s reasoning process, ultimately degrading performance.
Table 7: Performance of CFG under different starting steps.
Starting step Success Rate (%)
0 62.7
2 63.8
4 64.8
8 64.5
Appendix G Failure Cases of Track2Mask
From the Track2Mask pipeline presented in Appendix Fig.7, the failure cases fall into two categories: a) Object Detection Failure—the off-the-shelf open-vocabulary detector (i.e., GDINO) fails to localize objects in the initial observation; b) Incomplete Object Masking—the target objects along the trajectory are partially masked, as shown in Fig. 9. According to Equation 2, when the first failure case occurs, PCD’s prediction becomes equivalent to the original baseline result.
We conduct a ablation study to assess the impact of the incomplete masking failure cases in Table 8, where $\beta$ indicates the ratio of masked pixels manually excluded. As can be seen, PCD’s performance progressively declines as $\beta$ increases, until it approaches the baseline performance when $\beta$ reaches 60%. Nevertheless, both kinds of failure cases are exceptionally rare in our experiments, owing to the stability and effectiveness of the employed off-the-shelf models, i.e., GDINO for object detection and SAM v2 for object tracking and segmentation.
Figure 9: Failure cases of the Track2Mask module.
Table 8: Performance of PCD under different ratios ($\beta$) of masked pixels manually excluded. The task is “pick coke can”.
Model Baseline +PCD ($\beta=0$) $\beta=0.2$ $\beta=0.4$ $\beta=0.6$
OpenVLA 25 40 32 28 27
$\pi_{0}$ 84 88 88 85 84
Appendix H Applicability to More Complex Tasks
In our experiments on multi-objects tasks, PCD masks all task-relevant objects simultaneously. Here, we perform a new ablation study using the "Move Near" task to validate this strategy. Specifically, for task “Move A Near B”, we compare the results of i) masking only object A, ii) masking only object B, iii) masking A first, followed by masking B upon a successful grasp of A; and iv) Masking both A and B. The results are shown in Table 9, where the baseline model is OpenVLA. As can be observed, masking all task specific objects simultaneously yields the best performance.
Table 9: Ablation study on masking strategies for multi-object tasks.
Model Task Mask A Mask B Mask A then B Mask A and B
OpenVLA Move Near 60 50 57 62
$\pi_{0}$ Carrot Plate 53 62 56 67
Our PCD method has been evaluated on the long-horizon task "Apple Drawer" in the SIMPLER environment. As illustrated in Appendix Table 3, the official language instruction of the task is composed of two sequential sub-tasks: "open the top drawer" and "place the apple into the top drawer". Our strategy is to mask objects based on the instruction of the current sub-task. For sub-task 1, we mask only the "top drawer"; for sub-task 2, we mask both the "apple" and the "top drawer". In other words, PCD holds potential for long-horizon tasks by leveraging an LLM-based planner with CoT reasoning to decompose abstract commands into concrete sub-tasks.
Appendix I Performance Gaps among Different Object Annotation Strategies
The observed performance gaps among different object annotation strategies in Table 1 arises from the distinct "preferences" inherent in each approach. For instance, prompts from humans versus those from object detection models exhibit different understandings of object presence. As illustrated in Fig.10, in the case of a "drawer", a human annotator tends to point to or box a specific drawer within a cabinet, whereas an object detection model typically bounds the entire cabinet containing all drawers. This leads to significant performance disparities on "drawer"-related tasks. We argue that no single object annotation strategy is universally optimal; rather, the most effective approach is contingent upon the specific task.
Figure 10: Object masking results of Track2Mask using different object annotation strageties.
Appendix J Robustness under Various Magnitudes of Distractors
Here, we investigate the robustness of the PCD method under various magnitudes of distractors by incrementally increasing the number of distractors in the image background of the "pick coke can" task. The obtained results confirm that PCD maintains superior performance and exhibits greater robustness against these disturbances compared to the baseline.
Model Original Few distractors Many distractors
OpenVLA 25 24 14
+PCD 40 39 36
Table 10: Robustness of PCD to various magnitudes of distractors presented in Fig.11.
Figure 11: Various magnitudes of distractors.
Appendix K Performance in Multi-perspective Scenarios
In this part, we conduct an experiment to investigate PCD’s effectiveness in multi-perspective scenarios. Specifically, we leverage the miniVLA+VQ h8+Wrist [Belkhale and Sadigh, 2024] policy—which incorporates both third-person perspective and wrist images as input—as our baseline. The achieved results on five randomly sampled tasks from the LIBERO-90 [Liu et al., 2023a] benchmark are shown in Table 11, demonstrating that PCD consistently improves the baseline policy across all five tasks.
LIBER-90 Task miniVLA+VQ h8+Wrist +PCD
put the butter at the front in the top drawer of the cabinet and close it 54 72
put the black bowl on top of the cabinet 98 100
pick up the ketchup and put it in the basket 72 84
put the white mug on the plate 86 88
pick up the book and place it in the front compartment of the caddy 46 54
Average 71.2 79.6
Table 11: Performance of PCD on five LIBER-90 Tasks.
Appendix L LLMs Usage Statement